# An Introductory Survey to Autoencoder-based Deep Clustering - Sandboxes for Combining Clustering with Deep Learning

### v1

@@ 1. Introduction
There are two common ways to distinguish DC approaches: the optimization strategy (Nutakki et al., 2019; Zhou et al., 2025) and the type of neural network used (Ren et al., 2024; Min et al., 2018). We follow the taxonomy given in (Zhou et al., 2025) and divide the optimization strategies of DC methods into sequential (Section 3.1), alternating (Section 3.2), simultaneous (Section 3.3), and generative. Note that in the literature, sometimes other names for these groups are used, e.g., sequential = multi-stage = separate; alternating = iterative = closed-loop; simultaneous = joint. Sequential DC completely separates the clustering task from the representation learning task. Here, the NN is only used to initially learn a (clustering-independent) representation of the data. Alternating DC iteratively improves the representation and the clustering result, with the intermediate results influencing each other, while simultaneous methods optimize both objectives in parallel. Generative approaches, such as VaDE (Jiang et al., 2017) or ClusterGAN (Mukherjee et al., 2019), learn the parameters of a distribution and sample from that distribution to get a representation of the data, which can then be used for clustering.
@@
Focus of this survey: In contrast to more general DC surveys, e.g., (Zhou et al., 2025; Ren et al., 2024; Lu et al., 2024; Min et al., 2018; Huang et al., 2024), which provide a broad overview of the entire DC landscape, we discuss key concepts of AE-based applications in-depth (highlighted in green in Figure 1). We show the essential components of these concepts and give a detailed explanation of the underlying loss functions. We focus on AEs as they offer a perfect sandbox for developing novel DC methods for several reasons:
  • AEs can learn non-linear representations, revealing patterns that cannot be captured by linear methods, like Principal Component Analysis (Ding and He, 2004).
  • AEs are not limited to specific data types and require no additional information regarding the data, e.g., invariances.
@@ 3.1.2. Deep Embedding Network (DEN)
(6) $\mathcal{L}_{s}(\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{x_{i}\in\mathcal{B}}\sum_{g=1}^{k}\lambda\sqrt{n_{g}}{||z_{i}^{g}||}_{2}~{}$
@@ 3.1.4. Deep Density-based Image Clustering (DDC)
A negative aspect of N2D and DDC is that one cannot use the learned model to trivially predict the label of newly obtained data. The problem is that the underlying manifold learning algorithms (t-SNE, UMAP) do not learn a function $f(\cdot)$ to transform the data. Therefore, there is no straightforward way to embed new samples into the manifold where the labels can be obtained.
@@ 3.2. Alternating Deep Clustering
where $\lambda_{1}$ and $\lambda_{2}$ are introduced to weight the two losses and $C$ is the current clustering result (often consisting of cluster labels $Y$ and cluster centers $\mu$). Figure 3 visualizes the concept of obtaining a clustering-friendly embedding. Compared to Figure 2, one can see that the introduction of the clustering loss $\mathcal{L}_{clust}(\mathcal{B},C)$ leads to more compact groups of samples in the embedding of the AE.

### v2

@@ 1. Introduction
There are two common ways to distinguish DC approaches: the optimization strategy (Nutakki et al., 2019; Zhou et al., 2025) and the type of neural network used (Ren et al., 2024; Min et al., 2018; Wei et al., 2024). We follow the taxonomy given in (Zhou et al., 2025) and divide the optimization strategies of DC methods into sequential (Section 3.1), alternating (Section 3.2), simultaneous (Section 3.3), and generative. Note that in the literature, sometimes other names for these groups are used, e.g., sequential = multi-stage = separate; alternating = iterative = closed-loop; simultaneous = joint. Sequential DC completely separates the clustering task from the representation learning task. Here, the NN is only used to initially learn a (clustering-independent) representation of the data. Alternating DC iteratively improves the representation and the clustering result, with the intermediate results influencing each other, while simultaneous methods optimize both objectives in parallel. Generative approaches, such as VaDE (Jiang et al., 2017) or ClusterGAN (Mukherjee et al., 2019), learn the parameters of a distribution and sample from that distribution to get a representation of the data, which can then be used for clustering.
@@
Focus of this survey: In contrast to more general DC surveys, e.g., (Zhou et al., 2025; Ren et al., 2024; Min et al., 2018; Wei et al., 2024), which provide a broad overview of the entire DC landscape, or those focusing on image data (Huang et al., 2024) or prior knowledge (Lu et al., 2024), we discuss key concepts of AE-based applications in-depth (highlighted in green in Figure 1). We show the essential components of these concepts and give a detailed explanation of the underlying loss functions. We focus on AEs as they offer a perfect sandbox for developing novel DC methods for several reasons:
  • AEs can learn non-linear representations, revealing patterns that cannot be captured by linear methods, like Principal Component Analysis (PCA) (Pearson, 1901).
  • AEs are not limited to specific data types and do not require additional information regarding the data, e.g., invariances.
@@ 3.1.2. Deep Embedding Network (DEN)
(6) $\mathcal{L}_{s}(\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{x_{i}\in\mathcal{B}}\sum_{g=1}^{k}\lambda\sqrt{n_{g}}{||z_{i}^{g}||}_{2}.$
@@ 3.1.4. Deep Density-based Image Clustering (DDC)
A negative aspect of N2D and DDC is that one cannot use the learned model to trivially predict the labels of newly obtained data. The problem is that the underlying manifold learning algorithms (t-SNE, UMAP) do not learn a function $f(\cdot)$ to transform the data. Therefore, there is no straightforward way to embed new samples into the manifold where the labels can be obtained.
@@ 3.2. Alternating Deep Clustering
where $\lambda_{1}$ and $\lambda_{2}$ are introduced to weight the two losses and $C$ is the current clustering result (often consisting of cluster labels $Y$ and cluster centers $\mu$). For simplification, $\lambda_{1}=1$ is often chosen. Figure 3 visualizes the concept of obtaining a clustering-friendly embedding. Compared to Figure 2, one can see that the introduction of the clustering loss $\mathcal{L}_{clust}$ leads to more compact groups of samples in the embedding of the AE.
@@ 7. Conclusion
Appendix A Used Parameters
As highlighted in Section 5.1, selecting appropriate hyperparameters for (AE-based) deep clustering (DC) methods can be particularly challenging. To provide some intuition on this issue, we present the hyperparameter settings used by the DC algorithms covered in this survey, as stated in the respective papers. Specifically, Table 3 lists the hyperparameters for sequential methods, Table 4 for alternating methods, and Table 5 for simultaneous methods. Note that another important factor in reproducing certain results is the pre-processing of the data, e.g., by applying standardization or normalization.
Table 3. Parameterization of the discussed sequential deep clustering algorithms as described in the corresponding publications. * indicates that the parameter is automatically determined and $\dagger$ indicates that no information is given. If values are specified for a specific data set, this is indicated in brackets.
AE+$k$-Means DEN (Huang et al., 2014) N2D (McConville et al., 2020) DDC (Ren et al., 2020)
batch size $\beta$ Depending on the general setting - see pre-training parameters of alternating and simultaneous methods $|\mathcal{X}|$dagger $|\mathcal{X}|$
[…]
number of epochs $\epsilon$dagger $1000$dagger
optimization parameters $\phi$ Initial layer-wise training optimizer =absent=$\dagger$= † $lr=\dagger$ optimizer = ADAM $lr=\dagger$ optimizer = $\dagger$ $lr=\dagger$
clustering parameters $\omega$ $\alpha=250$ (COIL20) $\alpha=0.5$ (YaleB) $\alpha=0$ (PIE) $\beta=0$ (COIL20/PIE) $\beta=1$ (YaleB) $n_{g}=\frac{m}{k}$ $J=\dagger$ $\lambda=\dagger$ $t=\dagger$ $k$-Means with $100$ restarts # neighbors for ISOMAP = 5 # neighbors for UMAP = 20 perplexity for t-SNE = $\dagger$ # output features = $k$ EM algorithm perplexity for t-SNE = $\dagger$ # output features = 2 $ratio=0.1$ (for the DPC extension)
self-supervised loss function $l$ see Eq. 4, where $\mathcal{L}_{rec}$ is MSE (Eq. 2) $\dagger$ (Probably MSE (Eq. 2)) MSE (Eq. 2)
Table 4. Parameterization of the discussed alternating deep clustering algorithms as described in the corresponding publications. * indicates that the parameter is automatically determined and $\dagger$ indicates that no information is given. If values are specified for a specific data set, this is indicated in brackets.
AEC (Song et al., 2013) DCN (Yang et al., 2017) ACe/DeC (Miklautz et al., 2021) DipEncoder (Leiber et al., 2022)
batch size $\beta$ $|\mathcal{X}|$ $0.01|\mathcal{X}|$ $256$ $256$ for pre-training $25\cdot k$ for clustering
autoencoder parameters $\theta$ Encoder = $d$10001000100010002502502502505050505010101010 Decoder = mirrored encoder Encoder = $d$10010010010050505050101010102222 (Synthetic data) Encoder = $d$200020002000200010001000100010001000100010001000-100010001000100050505050 (RCV1) Encoder = $d$25025025025010010010010020202020 (20Newsgroup) Encoder = $d$$500500$-200020002000200010101010 (MNIST) Encoder = $d$5050505020202020-$5$ (MNIST-Processed) Encoder = $d$$1616$-$10$ (Pendigits) Decoder = mirrored encoder Encoder = $d$500500500500500500500500200020002000200010101010 Decoder = mirrored encoder Encoder = $d$500500500500500500500500200020002000200010101010 Decoder = mirrored encoder
number of pre-train epochs $\epsilon_{1}$dagger $50$ (RCV1/MNIST/Pendigits) $10$ (20Newsgroup/MNIST-Processed) $150000$ iterations $\rightarrow\epsilon_{1}=\frac{\beta\cdot 150000}{N}$ $100$
number of clustering epochs $\epsilon_{2}$ $50$ $50$ $100000$ iterations $\rightarrow\epsilon_{2}=\frac{\beta\cdot 100000}{N}$ $100$
pre-train optimization parameters $\phi_{1}$dagger Initial layer-wise training optimizer = SGD momentum $=0.9$ $lr=0.01$ Initial layer-wise training optimizer = ADAM $lr=0.001$ $\beta_{1}=0.9$ $\beta_{2}=0.99$ dropout rate = 0.2 noise level = 0.2 weight decay = 0 optimizer = ADAM $lr=0.001$
clustering optimization parameters $\phi_{2}$ optimizer = SGD $lr=\dagger$ optimizer = SGD momentum $=0.9$ $lr=0.05$ (RCV1/MNIST) $lr=0.001$ (20Newsgroup) $lr=0.01$ (MNIST-Processed/Pendigits) optimizer = ADAM $lr=0.001$ scheduler = $\frac{lr}{2}$ every 20000 iterations $\beta_{1}=0.9$ $\beta_{2}=0.99$ optimizer = ADAM $lr=0.0001$
initial clustering parameters $\omega_{1}$ Random $L$ $k$-Means 20 repetitions: random initial rotation $V$ random initial hard feature assignments $\beta$ $k$-Means within the clustering space $k$-Means
deep clustering parameters $\omega_{2}$ None None None None
self-supervised loss function $l$ MSE (Eq. 2) MSE (Eq. 2) MSE (Eq. 2) for pre-training extended MSE–including $V$–for clustering MSE (Eq. 2)
weight for the reconstruction loss $\lambda_{1}$ $1$ $1$ $1$ (based on $\mathcal{L}_{rec}$ of first batch)
weight for the clustering loss $\lambda_{2}$ $0.1$ (MNIST) $0.6$ (USPS/YaleB) $0.1$ (RCV1/MNIST-Processed) $10$ (20Newsgroup) $1$ or $0.05$ (MNIST) $0.5$ (Pendigits) $1$ (robust w.r.t. different values) $1$
Table 5. Parameterization of the discussed simultaneous deep clustering algorithms as described in the corresponding publications. * indicates that the parameter is automatically determined and $\dagger$ indicates that no information is given. If values are specified for a specific data set, this is indicated in brackets.
DEC (Xie et al., 2016) IDEC (Guo et al., 2017a) DCEC (Guo et al., 2017b) DKM (Fard et al., 2020)
batch size $\beta$ $256$ $256$dagger $256$
autoencoder parameters $\theta$ Encoder = $d$500500500500500500500500200020002000200010101010 Decoder = mirrored encoder Encoder = $d$500500500500500500500500200020002000200010101010 Decoder = mirrored encoder Encoder = $\text{conv}_{32}^{5}-\text{conv}_{64}^{5}-\text{conv}_{128}^{3}-\text{FC}_{m=10}$ ($\text{conv}_{a}^{b}\triangleq$ conv layer with $a$ filters and $b\times b$ kernel) Decoder = mirrored encoder Encoder = $d$-$500$-$500$- $2000$-$k$ Decoder = mirrored encoder
number of pre-train epochs $\epsilon_{1}$ $300000$ iterations $\rightarrow\epsilon_{1}=\frac{\beta\cdot 300000}{N}$ $300000$ iterations $\rightarrow\epsilon_{1}=\frac{\beta\cdot 300000}{N}$ $200$ $50$ (00 if annealing $\alpha$)
number of clustering epochs $\epsilon_{2}$ (controlled by $tol$) * (controlled by $\delta$) * (controlled by $\delta$) $100$ ($200$ if annealing $\alpha$)
pre-train optimization parameters $\phi_{1}$ Initial layer-wise training optimizer = SGD $lr=0.1$ scheduler = $\frac{lr}{10}$ every 20000 iterations dropout rate = 0.2 weight decay = 0 Initial layer-wise training optimizer = SGD $lr=0.1$ scheduler = $\frac{lr}{10}$ every 20000 iterations dropout rate = 0.2 weight decay = 0 optimizer = ADAM $lr=\dagger$dagger
clustering optimization parameters $\phi_{2}$ optimizer = SGD $lr=0.01$ For MNIST: optimizer = ADAM $lr=0.001$ $\beta_{1}=0.9$ $\beta_{2}=0.999$ For USPS/Reuters10k: optimizer = SGD $lr=0.1$ momentum $=0.99$dagger optimizer = ADAM $lr=0.001$ $\beta_{1}=0.9$ $\beta_{2}=0.999$
initial clustering parameters $\omega_{1}$ $k$-Means with 20 restarts $k$-Means with 20 restarts $k$-Means $k$-Means
deep clustering parameters $\omega_{2}$ $\alpha=1$ $tol=0.1\%$ (stopping criterion) $\alpha=1$ $\delta=0.1\%$ (convergence threshold) $T=140$ (MNIST) $T=30$ (USPS) $T=3$ (Reuters10k) $\alpha=1$ $\delta=0.1\%$ (convergence threshold) $T=140$ Version 1 (constant $\alpha$): $\alpha=1000$ Version 2 (annealing $\alpha$): $\alpha_{1}=0.1,\alpha_{i+1}=\alpha_{i}2^{-\log(i)^{2}}$
self-supervised loss function $l$ MSE (Eq. 2) MSE (Eq. 2) MSE (Eq. 2) MSE (Eq. 2)
weight for the reconstruction loss $\lambda_{1}$ 00 $1$ $1$ $1$
weight for the clustering loss $\lambda_{2}$ $1$ $0.1$ $0.1$ Version 1 (constant $\alpha$): $1$ (MNIST/USPS) $0.1$ (20Newsgroup) $0.01$ (RCV1) Version 2 (annealing $\alpha$): $0.1$ (MNIST/USPS) $0.0001$ (20Newsgroup/RCV1)